# A Data Generation Perspective to the Mechanism of In-Context Learning

### v2

@@ 3 Skill Learning
The skill learning ability considers that the LLM can learn a new data generation function according to the in-context examples on the fly. For the interpretability of skill learning, we take the research line of function learning  [^3] We refer to algorithm learning as function learning with an emphasis on the approximated functions by algorithms and, in this way, it is easier to analyze ICL. for discussion. Specifically, pre-training is considered as a process to learn a class of functions that can fit the pre-training corpora, and the ICL inference is to learn a new data generation function to best fit ICL demonstrations.
The discussions on skill learning are organized as follows. In Section 3.1, we first provide a clear description of the function learning framework and illustrate its benefits and drawbacks. In Section 3.2, we investigate (1) the function fitting ability of ICL by examining whether ICL can learn new functions from a certain function class; and (2) the generalization ability of ICL. In Section 3.3, we explain the internal mechanisms why ICL can implement different learning algorithms.
@@ 3.1 The Function Learning Framework
The function learning framework is a prototype setting defining a new learning task. Denoting $\mathbf{x}\sim\mathcal{P}_{\mathcal{X}},\mathbf{x}\in\mathbb{R}^{d}$ where $\mathcal{P}_{\mathcal{X}}$ is a distribution, a function class $\mathcal{F}$ where for each $f\in\mathcal{F},f:\mathbb{R}^{d}\to\mathbb{R}$. Given a sequence ($\mathbf{x}_{1},\cdots\mathbf{x}_{i}$) ($i>1$) sampled from $\mathcal{P}_{\mathcal{X}}$ sequentially, and a sampled function $f\sim\mathcal{F}$, the learning objective aims to correctly predict $f(x_{i})$ based on the sequence $(\mathbf{x}_{1},f(\mathbf{x}_{1}),\cdots,\mathbf{x}_{i-1},f(\mathbf{x}_{i-1}),\mathbf{x}_{i})$ with both in-context examples and the query input $\mathbf{x}_{i}$.
[…]
@@
where $\mathcal{L}$ is the loss function. $T_{\mathbf{\omega}}$ denotes the transformer model, $\mathbf{\omega}$ is the parameter of the transformer.
@@
Notably, the model is pre-trained on an ICL objective rather than the original next-token prediction one. The function learning framework enables us to: (1) arbitrarily generate data with desired properties from the pre-defined function class $\mathcal{F}$; (2) clearly examine the function-fitting ability and the generalization of the skill learning in the ICL; and (3) utilize well-developed statistical learning theory to understand the ICL.
However, existing research on function learning is not sophisticated as it can not tell the characteristics of the learned function classes. For instance, some studies claim that ICL inference learns a gradient descent algorithm, but Fu et al. (2023a) shows a high-order version of the gradient descent algorithm. Besides, the function classes are learned implicitly during pre-training and ICL inference. Previous works prove this simulation by approximating the target algorithm, e.g., ridge regression, with a small training error gap between the ridge regression and self-attention networks (Bai et al., 2023; Li et al., 2023b). The simulation accuracy is not well discussed or compared. Moreover, this line of research has to consider both pre-training and ICL inference, and, due to the computational resource limitation, most works utilize transformers with less than 6 layers. Therefore, conclusions may not be generalizable to larger model scales.
3.2 Function-fitting and Generalization of ICL
In this subsection, we investigate the function fitting and the generalization behavior of the ICL. function fitting indicates to what extent transformers can learn unseen data generation functions in context based on the pre-training generation functions within the same class. Generalization, on the other hand, measures how well transformers can learn an unseen function best in the in-context demonstrations, and the in-context function is from one of the pre-training function classes, e.g. linear regression and decision tree.
@@ 3.2 Function-fitting and Generalization of ICL
To explore the function fitting, previous works validate if ICL inference can learn a similar function as the pre-training data generation function. Raventos et al. (2023) leverage separated data generation functions, sampled from a linear function class, for pre-training data and in-context demonstrations. With a smaller class of linear functions, transformers do not show the skill learning ability but the skill recognition one, and are a Bayesian optimal estimator, for the data generation function of in-context demonstrations, given the pre-training functions as a prior. Provided with a larger class of linear functions, all unseen functions in the linear function class can be implemented by skill learning, achieving a performance matching the optimal least squares estimator. Wu et al. (2023) further provides a theoretical explanation to support the above empirical observations.
@@
Beyond the linear function class, Garg et al. (2022) observes that the ICL is expressive enough to implement more complicated functions when pre-training functions are also from the same function class, including sparse linear functions, two-layer neural networks, and decision trees. Bai et al. (2023); Fu et al. (2023b) propose theoretical explanations with a generalization bound between the prediction error of the toy transformer model and that of the target function, e.g. decision tree. However, there are still two essential questions unsolved: (1) Why the transformers suddenly obtain the skill learning ability with significant performance increase once the number of pre-training data generation functions reaches a certain threshold? (2) Why the learned data generation function of ICL demonstrations is from the same class as the pre-training data generation function?
@@
The generalization of ICL is validated by comparing the ground-truth data generation function of in-context demonstrations and the predicted one. Assuming the pre-training data generation functions cover decision tree and linear function, the ground-truth data generation function of ICL demonstrations is a linear function. The ICL generalization is strong if and only if the predicted function of ICL demonstrations is a linear function.
@@
Bai et al. (2023); Ahuja et al. (2023) indicate that the transformer can implement the Bayesian optimal selection which selects the best-fitting function, with the minimum description length, from all the function classes that appeared in the pre-training stage. Such Bayesian optimal selection helps a transformer pre-trained with multiple function classes reach comparable ICL performance as one pre-trained with the ground-truth function class. Notably, such Bayesian optimal on the synthetic dataset may not extend to the real-world More surprisingly, Yadlowsky et al. (2023) observes even if each pre-training instance is not generated from one individual function class. Instead, each instance is generated from a weighted sum of multiple functions from different function classes, e.g., $0.7f_{1}(x)+0.3f_{2}(x)$ where $f_{1}$ and $f_{2}$ are from different function classes. The same conclusion holds. Nonetheless, skill learning fails if the ground-truth function is out of the pre-training function class, then the transformer only demonstrates skill recognition ability (Yadlowsky et al., 2023).
@@
Summary: Skill learning ability emerges if the function class of potential pre-training data generation functions is sufficiently large. For in-context demonstrations, ICL can learn a function that lies in the same function class of the pre-training data. The ICL would implement a Bayesian optimal selection to select the best-fitting function, aligned with ICL demonstrations generate function, from pre-training function classes.
@@ 3.3 The Internal Mechanisms of ICL
In this subsection, we further explore the mechanism of ICL by answering why it can learn an unseen function in context. Notably, there are two common assumptions in existing works: (1) The data generation functions for both pre-training data and in-context demonstrations are linear. (2) The toy transformer model is linearized by removing feed-forward layers and the softmax activation function in the attention layer. Ahn et al. (2023b) proposes indirect evidence that the linearized simplification does not hurt the precision of the vanilla transformer architecture, illustrating that the training dynamic of the linearized version is similar to that of the vanilla version.
@@
Previous works analogize ICL to meta-learning (Finn et al., 2017). The pre-training stage corresponds to the outer-loop optimization, and the ICL inference stage is an instance of the inner-loop optimization, implementing fast adaptation on new novel tasks. Rather than a real inner gradient update, ICL inference mimics gradient update via a forward process with in-context demonstrations (Hubinger et al., 2019; von Oswald et al., 2023). Based on the dual view that a linear neural layer updated with gradient is equivalent to a linear attention forward process (Irie et al., 2022)Dai et al. (2022) proves the mathematical equivalence, illustrating the implicit gradient descent implementation with a linear attention. However, such an analogy is only limited to mathematical equivalence, it is not a straightforward answer to why ICL can learn a function since such an analogy overlooks all practical details on the gradient descent, including the choice of the learning objective, the pre-training weight, and the training data distribution.
@@
Following-up research focuses on how the ICL forward process can implement different functions, given suitable assumptions on the transformer’s weights. Von Oswald et al. (2023) first demonstrate that the ICL on the single-layer transformer can implement one-step gradient descent with a linear regression objective. Bai et al. (2023) further show the ICL inference can implement ridge regression, least square, lasso, and even gradient descent on a two-layer Neural Network. Nonetheless, those strong assumptions about the attention weights may be not practically reasonable. For instance, Von Oswald et al. (2023) construct the key, query, value matrices $W_{K},W_{Q},W_{V}$ with $W_{K}=W_{Q}=\left(\begin{array}[]{cc}I_{x}&0\\ 0&0\end{array}\right),W_{V}=\left(\begin{array}[]{cc}0&0\\ W_{0}&-I_{y}\end{array}\right)$, where $I_{x}$ and $I_{y}$ are two different identity matrices and $W_{0}$ is the initialized parameters of the transformer model. Nonetheless, it is unclear why a pre-trained transformer would have such type of weights, and it has been reported not easily achieved in practice (Shen et al., 2023).
@@
To extend the construction of attention weights aforementioned into a more practical setting, Zhang et al. (2023a); Mahankali et al. (2023); Ahn et al. (2023a) showcase the converged weights after the pre-training stage can simulate some learning algorithms during the ICL inference. However, the converged weights may not be precise enough since they are determined based on the loss landscape or gradient flow analysis. Typically, the ICL on the one-layer linear transformer can implement gradient descent (Von Oswald et al., 2023) or preconditioned gradient descent algorithm (Ahn et al., 2023a) given a linear regression objective. Given a two-layer transformer, ICL can implement a gradient descent with adaptive step size and special sparsity regularization (Ahn et al., 2023a). Moreover, Ahn et al. (2023a); Von Oswald et al. (2023) reveal that multiple-layered transformers can implement a GD++ algorithm. For larger-scale transformers, Akyürek et al. (2022) empirically illustrates that, instead of performing GD, large-scale transformers show emergent ability directly approximating the closed-form solution of ridge-regression, while there is still a gap on why this ability emerges as the model-scale increases.
@@
Beyond the linear activation for attention heads, more recent research takes the softmax activation function into consideration. Von Oswald et al. (2023) demonstrate there exists a transformer that performs GD to solve more complicated nonlinear regression tasks. Li et al. (2023a); Ren & Liu (2023) identify the nonlinear regression task as the softmax regression and contrastive learning objective, respectively. Cheng et al. (2023) further take non-linear data generation functions into consideration, elucidating a transformer can implement gradient descent and converge to the Bayes optimal predictor. Guo et al. (2023) further studies a more challenging but practical setting of representation learning, in which predictions depend on inputs through the MLP. The theoretical evidence indicates that the ICL inference can implement ridge regression in context with the input of neural representations.
Summary: Existing research shows the mathematical equivalence between gradient descent and toy transformer models, working as proof that transformers implicitly implement gradient descent in context. However, this is supportive evidence for the function learning framework while there are no insights about how to select in-context demonstrations.
4 Insights on the Bayesian Inference and the Function Learning Framework
@@
The skill recognition and skill learning abilities are comprehensively discussed with the example statistical frameworks of the Bayesian inference and function learning in Section 2 and 3, respectively. Instead of explaining one ability with one specific statistical framework, we exhibit the potential that both frameworks can be easily utilized to understand the mechanism of both abilities. Such extension enables the future mechanism analysis to select the suitable analysis framework, by referring to their strengths and weaknesses. The function learning framework provides an elegant description of the data generation process with more comprehensive conclusions. However, it is over-simplified with an unclear relevance to the real-world scenario. The Bayesian inference framework provides a more concrete and detailed description of the data generation process through an HMM model, e.g., the delimiter is taken into consideration, while the theoretical analysis on the role of delimiters is hard since it requires several assumptions over statistical modeling.
@@
The function learning framework can be easily extended to understand skill recognition by simply replacing the data generation function from a mixture of HMMs with a mixture of linear functions. In this section, we focus on how to utilize the Bayesian inference framework to model the mechanism of skill learning. We first show that the original function learning framework for the skill learning ability also implements an implicit Bayesian optimal selection in Section 4.1. We then extend the Bayesian inference framework to learn new in-context data generate functions in Section 4.2, therefore the Bayesian inference framework can also be a technical solution for skill learning.
4.1 Bayesian Selection in the Function Learning Framework

### v3

@@ 5 Skill Learning
Through the skill learning ability, LLMs can inference a new data generation function which has not been seen during pre-training. The function learning framework [^2] We refer to algorithm learning as function learning with an emphasis on the approximated functions by algorithms and, in this way, it is easier to analyze ICL. is utilized to interpret the skill learning ability. Specifically, pre-training is considered as a process to learn a class of functions that can fit the pre-training corpora, and the ICL inference is to learn a new data generation function via fitting the ICL demonstrations.
Discussions on the skill learning ability are organized as follows. In Section 5.1, we first provide a clear description of the function learning framework and illustrate its benefits and drawbacks. In Section 5.2, we investigate: (1) whether LLMs can learn new functions in context, and (2) if so, whether the learned functions can effectively generalize to test samples.
In Section 5.3 illustrates ICL can implement different learning algorithms, e.g., gradient descent. More discussions on the robustness of ICL can be found in Appendix F.
@@ 5.1 The Function Learning Framework
Previous research reformulates the pre-training objective of next-token prediction into an input-label mapping objective during the ICL inference stage. One limitation of the function learning framework is that it has to pre-train the model from scratch as the pre-training objective is different from the next token prediction. Due to computational resource limitations, most works utilize transformers with less than 6 layers. These conclusions may not be generalizable to larger scale models. Garg et al. (2022) has been the only work to utilize a relative larger-scale model, reaching a similar scale as GPT-2.
@@
Denoting $\mathbf{x}\sim\mathcal{P}_{\mathcal{X}},\mathbf{x}\in\mathbb{R}^{d}$ where $\mathcal{P}_{\mathcal{X}}$ is a distribution, a function class $\mathcal{F}$ where for each $f\in\mathcal{F},f:\mathbb{R}^{d}\to\mathbb{R}$. Given a sequence ($\mathbf{x}_{1},\cdots\mathbf{x}_{i}$) ($i>1$) sampled from $\mathcal{P}_{\mathcal{X}}$ sequentially, and a sampled function $f\sim\mathcal{F}$, the learning objective aims to correctly predict $f(x_{i})$ based on the sequence $(\mathbf{x}_{1},f(\mathbf{x}_{1}),\cdots,\mathbf{x}_{i-1},f(\mathbf{x}_{i-1}),\mathbf{x}_{i})$ with both in-context examples and the query input $\mathbf{x}_{i}$.
$\underset{\begin{subarray}{c}\mathbf{x}_{1}\dots\mathbf{x}_{n}\sim\mathcal{P}_{\mathcal{X}}\\ f\sim\mathcal{F}\end{subarray}}{\mathbb{E}}\left[\sum_{i=2}^{n}\!\mathcal{L}\left(f\!\left(\mathbf{x}_{i}\right),T_{\omega}\left(\left[\mathbf{x}_{1},f\left(\mathbf{x}_{1}\right)\ldots\mathbf{x}_{i}\right]\right)\right)\right]$ (2)
@@
Eq. (2) describes the learning objective, where $\mathcal{L}$ is the loss function. $T_{\mathbf{\omega}}$ denotes the transformer model, $\mathbf{\omega}$ is the parameter of the transformer.
@@
Notably, the model is pre-trained on the above ICL objective instead of the original next-token prediction objective. The function learning framework enables us to: (1) arbitrarily generate data with desired properties from the pre-defined function class $\mathcal{F}$; (2) clearly examine the function-approximation ability and the generalization of skill learning in ICL; and (3) utilize well-developed statistical learning theory to understand ICL.
5.2 Function Approximation and Generalization of ICL
In this subsection, we investigate the function approximation and generalization behavior of ICL. Function approximation indicates to what extent transformers can approximate the ground-truth function underlying a given input, in the ICL inference stage. Generalization, on the other hand, measures the gap between the approximated function and the ground-truth data generation function. Notably, the function learning framework investigates ICL in the function space, rather than the token space.
@@ 5.2 Function Approximation and Generalization of ICL
To explore the function approximation ability, Raventos et al. (2023) leverages different linear functions to generate pre-training data and in-context demonstrations. When pre-training on a small set of linear functions, ICL acts as a Bayesian optimal estimator, illustrating the skill recognition ability Raventos et al. (2023). If enlarging the set of pre-training linear functions, ICL can act as an optimal least squares estimator with better function approximation, illustrating the skill learning ability Raventos et al. (2023). Wu et al. (2023a) provides a theoretical explanation to support the above empirical observations.
@@
Beyond the linear function class, Garg et al. (2022) observes that the ICL is expressive enough to approximate more complicated functions, including sparse linear functions, two-layer neural networks, and decision trees. The only requirement is that the same function class must be encountered during both pre-training and the ICL stage. Bai et al. (2023) and Fu et al. (2023a) establish a statistical task complexity bound for pre-training, supporting the empirical observations mentioned above. The findings suggest that skill learning can be achieved with a dimension-independent number of linear regression pre-training tasks. However, two essential questions remain unsolved: (1) Why do transformers suddenly obtain the skill learning ability with significant performance increase once the number of pre-training data generation functions reaches a certain threshold? (2) Why is the learned data generation function of ICL demonstrations from the same class as the pre-training data generation function?
@@
The generalization of ICL is validated by comparing the ground-truth data generation function of in-context demonstrations and the approximated one through ICL inference. A more complicated experimental setting is considered where pre-training involves data generation functions from multiple function classes simultaneously, rather than being restricted to a single function class, as in the above function approximation experiments. Assuming pre-training data generation functions cover decision trees and linear functions, the ground-truth data generation function of ICL demonstrations is a linear function. The ICL generalization is strong if and only if the predicted function of ICL demonstrations is a linear one.
@@
Bai et al. (2023); Ahuja et al. (2023); Vasudeva et al. (2024); Tripuraneni et al. (2023) indicate that transformers can achieve the Bayesian optimal selection, choosing the best-fitting function class with the minimum description length, from those function classes seen during the pre-training stage. Such Bayesian optimal selection helps a transformer pre-trained with multiple function classes reach comparable ICL performance as one pre-trained with only the ground-truth function class. Notably, such Bayesian optimal on the synthetic dataset may not fully explain all the experimental observations. Yadlowsky et al. (2023) generates each pre-training instance with functions from multiple function classes, e.g., $0.7f_{1}(x)+0.3f_{2}(x)$ where $f_{1}$ and $f_{2}$ are from different function classes. The ICL can still achieve Bayesian optimal selection, holding the same conclusion. Notably, the above works focus on the scenario where the ground-truth data function is within pre-training function classes. Skill learning fails if the ground-truth data function is out of the pre-training function class Yadlowsky et al. (2023); ICL degrades to skill recognition with Bayesian optimal estimator.
@@
In summary, skill learning emerges if the number of pre-training data generation functions is sufficiently large. ICL can learn a function that lies in the same function class of the pre-training data. Moreover, ICL would implement a Bayesian optimal selection to select the function best-fitting on ICL demonstrations, from pre-training function classes.
@@ 5.3 The Internal Mechanisms of ICL
In this subsection, we explore how ICL can learn an unseen function in context. Notably, there are two common assumptions generally utilized in existing works: (1) The data generation functions for both pre-training data and in-context demonstrations are linear. (2) The toy transformer model is linearized by removing feed-forward layers and the softmax activation function in the attention layer. This linearized simplification may generalize to the standard transformer, as Ahn et al. (2023b) illustrates that the training dynamic of the linearized version is similar to the standard transformer.
@@
Previous works analogize ICL to meta-learning Finn et al. (2017). The pre-training stage corresponds to the outer-loop optimization, and the ICL inference stage is an instance of the inner-loop optimization, implementing fast adaptation on new novel tasks. Rather than a real inner gradient update, ICL inference mimics gradient update via a forward process with in-context demonstrations Hubinger et al. (2019); von Oswald et al. (2023); Zheng et al. (2024).
Based on the dual view that the backward process on a linear neural layer is equivalent to the forward process on a linear attention layer, Irie et al. (2022); Dai et al. (2022) proves the mathematical equivalence, illustrating the implicit gradient descent implementation with a linear attention. However, such an analogy is only limited to mathematical equivalence. It remains unclear why ICL can learn a function since such an analogy overlooks many practical details, including the choice of the learning objective, pre-training weights, and the training data distribution Mahdavi et al. (2024).
@@
To address the gap between theoretical models and real-world implementation, the following works consider the construction of pre-training weights. Von Oswald et al. (2023) first demonstrate that ICL on the single-layer transformer can implement one-step gradient descent with a linear regression objective. Bai et al. (2023) further show that ICL inference can implement ridge regression, least square, lasso, and even gradient descent on a two-layer Neural Network. Nonetheless, those strong assumptions about the attention weights may be not practically reasonable. For instance, Von Oswald et al. (2023) construct the key, query, value matrices $W_{K},W_{Q},W_{V}$ with $W_{K}=W_{Q}=\left(\begin{array}[]{cc}I_{x}&0\\ 0&0\end{array}\right),W_{V}=\left(\begin{array}[]{cc}0&0\\ W_{0}&-I_{y}\end{array}\right)$, where $I_{x}$ and $I_{y}$ are two different identity matrices and $W_{0}$ is the initialized parameters of the transformer model. Nonetheless, it is unclear why a pre-trained transformer would have such type of weights, and it has been reported that this is not easily achieved in practice Shen et al. (2023).
@@
Instead of explicit attention weight construction, Zhang et al. (2023a); Mahankali et al. (2023); Ahn et al. (2023a) analyze the converged weights obtained after pre-training. Von Oswald et al. (2023) observes the ICL on the one-layer linear transformer can implement gradient descent or preconditioned gradient descent algorithm Ahn et al. (2023a) given a linear regression objective. Given a two-layer transformer, ICL can implement a gradient descent with adaptive step size and special sparsity regularization Ahn et al. (2023a). Moreover, Ahn et al. (2023a); Von Oswald et al. (2023) reveal that multiple-layered transformers can implement a GD++ algorithm. For larger-scale transformers, Akyürek et al. (2022) empirically illustrates that, instead of performing GD, large-scale transformers show emergent ability directly approximating the closed-form solution of ridge-regression, while there is still a gap on why this ability emerges as the model-scale increases.
@@
Beyond the linear activation for attention heads, recent researches take the softmax activation function into consideration. Von Oswald et al. (2023) demonstrates there exists a transformer that performs GD to solve more complicated nonlinear regression tasks. Li et al. (2023a); Ren and Liu (2023) identify the nonlinear regression task as the softmax regression and contrastive learning objective, respectively. Cheng et al. (2023) further takes non-linear data generation functions into consideration, elucidating a transformer can implement gradient descent and converge to the Bayes optimal predictor. Wibisono and Wang (2023) theoretically finds that the softmax can help to find the correct data pair from the unstructured data which the input-output pair is permuted. Guo et al. (2023); Zhang et al. (2024) further studies a more challenging but practical setting of representation learning, in which predictions depend on inputs through the MLP. The theoretical evidence in Guo et al. (2023) indicates that the ICL inference can implement ridge regression in context with the input of neural representations, while Collins et al. (2024) argue theoretically and empirically that ICL inference with a single self-attention head behaves like a Nadaraya-Watson kernel regressor and training the attention weights entails learning the appropriate neighborhood size and subspace for this regressor based on the Lipschitzness of the target functions.
Practical usage of mechanism analysis. The above section has indicated that ICL implements a gradient descent vector to achieve successful function learning. From a practical perspective, Todd et al. (2023); Liu et al. (2023b) find the existence of compressed task vectors [^3] Similar task vectors Hojel et al. (2024) can also be found in the computational vision domain. in transformers with specific functionality. More recently, Li et al. (2024) attempts to connect the gradient vector with the compressed task vector, utilizing inner and momentum optimization towards a better task vector. Success of the new optimized task vector can be found on multiple tasks.
6 Insights & Future Directions
In this section, we delve into key insights from the data mechanism perspective of ICL and identify open questions that remain to be addressed in this evolving field.
The uniformity of the two frameworks. The core idea from the data-generative perspective is to (1) construct a data generation function hypothesis with one specific statistical framework and (2) analyze the data generation capability of the LLM with ICL instances with a focus on either skill learning/recognition mechanism. The existing pipelines on skill recognition and skill learning abilities are comprehensively discussed with the statistical frameworks of the Bayesian inference and function learning in Section 4 and 5, respectively. However, most existing analysis follows one-to-one correspondence which explains one ability with one specific statistical framework, serving as a solution for skill learning.
@@ 6 Insights & Future Directions
Our new data generative perspective suggests the researcher find a suitable statistical framework as the starting point for analysis. We exhibit the potential that both frameworks can be easily utilized to understand the mechanism of both abilities. Such extension enables the future mechanism analysis to select the suitable analysis framework, by referring to their strengths and weaknesses. The function learning framework provides an elegant description of the data generation process with more comprehensive conclusions. However, it is over-simplified with an unclear relevance to the real-world scenario. The Bayesian inference framework provides a more concrete and detailed description of the data generation process through an HMM model, e.g., the delimiter is taken into consideration, while the theoretical analysis on the role of delimiters is hard since it requires several assumptions over statistical modeling.
@@
We provide a comprehensive discussion on extending one framework to the other statistical framework. The function learning framework can be easily extended to understand skill recognition by simply replacing the data generation function from a mixture of HMMs with linear functions. A comprehensive discussion on how to utilize the Bayesian inference framework to model the mechanism of skill learning in Appendix A. We first show that the original function learning framework for the skill learning ability also implements an implicit Bayesian optimal selection in Appendix A.1. We then extend the Bayesian inference framework to learn new in-context data generate functions in Appendix A.2. The Bayesian inference framework can also serve as a solution for skill learning.
The unique strengths and weaknesses of skill learning/recognition ability Considering the intricate interplay of both abilities on different tasks, we further illustrate the strengths and weaknesses inherent in each ability. Skill learning ability can obtain new knowledge from the in-context data, and even over-ride the pre-training knowledge. It provides an easy way to update the knowledge on the specific application without requiring computationally heavy fine-tuning. Such ability has been successfully utilized in different LLM applications, e.g. model editing with ICL Zheng et al. (2023). Nonetheless, the skill learning ability may fail as it can be easily distracted by irrelevant context Shi et al. (2023). Skill recognition ability is insensitive to the new in-context pattern leading to the failure on the specification-heavy task Peng et al. (2023) while it exhibits robustness to the incorrectness of label-demonstrations and other in-context noise Webson and Pavlick (2021). Based on the above discussion, we suggest a careful evaluation of LLMs about each ability and select a desired one for the downstream task.
Emergent Skill Composition Ability. We majorly focus on the skill recognition/learning ability in our paper. More recently, new skill composition ability is found on larger model with specialized ICL prompts like Chain-of-Thought (CoT) Wei et al. (2022b). The skill composition ability combines multiple data generation functions to create a more complicated data generation function. This ability, supported theoretically by Arora and Goyal (2023), shows that complex tasks can exhibit performance gains when decomposed skills improve linearly. More analyses on the effectiveness of skill composition ability can be found in Appendix C.
Application of Skills. After the LLM obtained the skill learning and skill recognition abilities during pre-training, we then investigate how the model utilizes both abilities for achieving satisfactory downstream task performance during the ICL inference stage. Overall, the behavior of the LLM is more consistent with the skill recognition mechanism on difficult tasks while observations aligned with skill learning are more common to see on easy tasks.
How the skill learning ability emerges during pre-training. The emergence of the skill learning ability can be partially attributed to the skewed rank-frequency distribution of pre-training corpora.  Chan et al. (2022a), and Reddy (2023) highlight the role of the induction head Olsson et al. (2022), a particular attention head which explicitly searches for a prior occurrence of the current token in-context and copying the suffix as predictions. Moreover, the function class-based analysis Raventos et al. (2023) illustrates that the transition from skill recognition to skill learning only happens given diverse enough tasks in pre-training corpora. It is interesting to explore how these factors collaboratively influence the emergence of skill learning.
Why does ICL only learn the data generation function that appeared during pre-training? In Section 5, we provide a comprehensive discussion on what function can be learned in context. Observations indicate that ICL can only learn the function within the pre-training data generation function class. Nonetheless, the causality of the pre-training data generation function to ICL remains unclear. Garg et al. (2022) proposes the research question as: Can we train a model to in-context learn a certain function class but overlooks the effect of the pre-training data generation function class. Once we have a certain clue about causality, we can leverage the skill-learning ability in a more controllable and safe manner.
Another line of research is to conduct analyses on more realistic scenarios. Recently, Chen et al. (2024) finds the parallel structures in pre-training data-pairs of phrases following similar templates in the same context window is the key to the emergence of the ICL capability. We conjecture that the underlying reason can be the formulation of the induction head with repeat patterns.
Data generation functions aligned with real-world scenarios. One major concern on the statistical framework is that the correspondence with real-world scenarios is unknown and overly simplified. Recently, Akyürek et al. (2024) proposes a new approach for generating data functions that are more aligned with real-world scenarios. The framework allows for more accurate simulations and testing of machine learning models by integrating domain-specific knowledge and constraints into the data generation process. This alignment enhances the applicability and reliability of existing conclusions to the real-world scenarios. We advocate for theoretical analyses focused on real-world data generation functions, moving beyond traditional statistical frameworks. More empirical analysis on skill learning and skill recognition abilities are illustrated in Appendix B.
Extending existing findings to other capabilities of LLMs. More ICL capabilities are observed except for classification tasks, e.g., step-by-step reasoning ability Wei et al. (2022b) for reasoning and self-correction Ganguli et al. (2023). A critical question is how we can extend the understanding frameworks introduced in this paper, particularly the data generation perspective, to more complicated LLMs’ capabilities. Some pioneering research has been done; Prystawski and Goodman (2023) extends the Bayesian inference framework to understand the effectiveness of the CoT prompt. Kadavath et al. (2022) focuses on the self-evaluation prompt showing that LLMs can accurately examine the correctness of their statements. We believe the introduced data generation perspective and two main understanding frameworks on ICL serve as the milestone to explore more intrinsic capabilities of LLMs.
7 Conclusion
In this study, we introduce a novel data generation perspective to understand the underlying mechanism driving the current success of ICL. We primarily focus on understanding the LLM’s ability of skill learning and skill recognition, and investigate whether ICL inference is capable of learning new data generation functions in context. Our work makes a step forward to enhancing our understanding of underlying mechanisms.
8 Limitations
In this paper, we provide a mechanism understanding of the ICL from a data generation perspective, We systematically consider the limitations from various perspectives such as fairness, security, harm to people, and so on, and we do not find any apparent social risk related to our work. However, there is a notable technical limitation in our study. The current statistical frameworks with controlled experimental settings may not fully capture complexities present in real-world scenarios. This gap between the theoretical framework and practical applications suggests that further research is needed to adapt and refine the mechanism analysis to align with real-world application.
Acknowledgement
We extend our gratitude to Dr. Liam Collins for his insightful feedback on this paper. Haitao Mao and Jiliang Tang are supported by the National Science Foundation under grant numbers CNS2321416, IIS2212032, IIS2212144, IOS2107215, DUE2234015, CNS2246050, DRL2405483 and IOS2035472, the Army Research Office under grant number W911NF-21-1-0198, Amazon Faculty Award, JP Morgan Faculty Award, Meta, Microsoft and SNAP.
Appendix A Insights on the Bayesian Inference and the Function Learning Framework
A.1 Bayesian Selection in the Function Learning Framework